### v1

# Feedback control of the Kuramoto model defined on uniform graphs I: Deterministic natural frequencies

Abstract.
We study feedback control of the Kuramoto model (KM) with uniformly spaced natural frequencies defined on uniform graphs which may be complete, random dense or random sparse, so that each node is desired to exhibit a constant rotational motion. We assume that the node number $n\geq 3$ and prove that in the controlled Kuramoto model (CKM) on complete graphs a synchronized solution that converges to the desired motion suffers a saddle-node bifurcation where a pair of stable and unstable ones born and that the corresponding solution is always asymptotically stable in the continuous limit (CL) if it exists. They are also proved to be the only synchronized solutions that are or can be stable in the CKM and CL, respectively, when the feedback gain is positive and the coupling constant is small enough for synchronized motions not to exist in the uncontrolled case. Moreover, we show based on the previous results that the asymptotically stable solution to the CL behaves as if it is an asymptotically stable one in the CKM when the node number is sufficiently large, even if the graphs are random dense or sparse, when the node number is sufficiently large. We demonstrate the theoretical results by numerical simulations for the CKM on complete simple, and uniform random dense and sparse graphs.
Key words and phrases:
Kuramoto model; feedback control; continuum limit; synchronization; bifurcation; stability
2020 Mathematics Subject Classification:
34C15; 34H05; 45J05; 34D06; 34C23; 37G10; 34D20; 45M10; 05C90
1. Introduction
We consider the Kuramoto model (KM) [17, 18] with natural frequencies defined on graphs $G_{n}$, $n\in\mathbb{N}$,
$\frac{\mathrm{d}}{\mathrm{d}t}u_{i}^{n}(t)=$ $~{}\omega_{i}^{n}+\frac{K}{n\alpha_{n}}\sum_{j=1}^{n}w_{ij}^{n}\sin\left(u_{j}^{n}(t)-u_{i}^{n}(t)\right),\quad i\in[n]:=\{1,2,\ldots,n\},$ (1.1)
where $n$ is the node number; $u_{i}^{n}(t)\in\mathbb{S}^{1}=\mathbb{R}/2\pi\mathbb{Z}$ and $\omega_{i}^{n}$, respectively, represent the phase and natural frequency of the oscillator at the node $i\in[n]$; $K>0$ is a coupling constant; and $\alpha_{n}$ is a scaling factor which is one if $G_{n}$ is dense. Here $G_{n}=\langle V(G_{n}),E(G_{n}),W(G_{n})\rangle$, $n\in\mathbb{N}$, represent a sequence of weighted graphs, where $V(G_{n})=[n]$ and $E(G_{n})$ are the sets of nodes and edges, respectively, and $W(G_{n})$ is an $n\times n$ weight matrix given by
$(W(G_{n}))_{ij}=\begin{cases}w_{ij}^{n}&\mbox{if $(i,j)\in E(G_{n})$};\\ 0&\rm{otherwise.}\end{cases}$
The edge set is expressed as
$E(G_{n})=\{(i,j)\in[n]^{2}\mid(W(G_{n}))_{ij}\neq 0\},$
where each edge is represented by an ordered pair of nodes $(i,j)$, which is also denoted by $j\to i$, and a loop is allowed. If $W(G_{n})$ is symmetric, then $G_{n}$ represents an undirected weighted graph and each edge is also denoted by $i\sim j$ instead of $j\to i$. When $G_{n}$ is a simple graph, $W(G_{n})$ is a matrix whose elements are $\{0,1\}$-valued. When $G_{n}$ is a random graph, $W(G_{n})$ is a random matrix. We say that the graph $G_{n}$ is dense if #$\#E(G_{n})/(\#V(G_{n}))^{2}>0$ as $n\rightarrow\infty$, and that it is sparse if #$\#E(G_{n})/(\#V(G_{n}))^{2}\rightarrow 0$ as $n\rightarrow\infty$.
The weight matrix $W(G_{n})$ is given as follows: Let $I=[0,1]$ and let $W\in L^{2}(I^{2})$ be a measurable function. If $G_{n}$, $n\in\mathbb{N}$, are deterministic dense graphs, then
$w_{ij}^{n}=\langle W\rangle_{ij}^{n}:=n^{2}\int_{I_{i}^{n}\times I_{j}^{n}}W(x,y)\mathrm{d}x\mathrm{d}y,$ (1.2)
where
$I_{i}^{n}:=\begin{cases}[(i-1)/n,i/n)&\mbox{for $i<n$};\\ [(n-1)/n,1]&\mbox{for $i=n$}.\end{cases}$
If $G_{n}$, $n\in\mathbb{N}$, are random dense graphs, then $w_{ij}^{n}=1$ with probability
$\mathbb{P}(j\rightarrow i)=\langle W\rangle_{ij}^{n},$ (1.3)
where the range of $W$ is contained in $I$. If $G_{n}$, $n\in\mathbb{N}$, are random sparse graphs, then $w_{ij}^{n}=1$ with probability
$\mathbb{P}(j\rightarrow i)=\alpha_{n}\langle\tilde{W}_{n}\rangle_{ij}^{n},\quad\tilde{W}(x,y):=\alpha_{n}^{-1}\wedge W(x,y),$ (1.4)
where $W$ is a nonnegative function, $\alpha_{n}=n^{-\gamma}$ with $\gamma\in(0,\frac{1}{2})$, and $a\wedge b=\min(a,b)$ for $a,b\in\mathbb{R}$. In such a situation, the function $W(x,y)$ is usually called a graphon[20].
Such coupled oscillators on complex networks have recently attracted much attention and have been extensively studied. They provide many mathematical models in various fields such as physics, chemistry, biology, social sciences and engineering. Among them, the KM is one of the most representative models and has been generalized in several directions. It has very frequently been subject to research, especially to discuss the synchronization phenomenon. See [31, 28, 1, 2, 5, 27, 29] for the reviews of vast literature on coupled oscillators in complex networks including the KM and its generalizations.
In the previous work [14], coupled oscillator networks such as (1.1) were studied and shown to be well approximated by the corresponding continuum limits (CLs), for instance, which is given by
$\frac{\partial}{\partial t}u(t,x)=\omega(x)+K\int_{I}W(x,y)\sin(u(t,y)-u(t,x))\mathrm{d}y,\quad x\in I,$ (1.5)
for (1.1) if the natural frequencies $\omega_{i}^{n}$, $i\in[n]$, are determined by the $L^{2}$ function $\omega(x)$ on $I$ as
[…]
We call $\omega(x)$ a frequency function and assume that it is not a constant function. More general cases in which networks of coupled oscillators are defined on multiple graphs which may be deterministic dense, random dense or random sparse were discussed in [14]. Some results on the stability of solutions to the coupled oscillators and CLs were also refined in [34]. Similar results for such networks which are defined on single graphs and have the same natural frequency were obtained earlier in [15, 21, 22, 24] although they are not applicable to (1.1) and (1.5). Such a CL was introduced for the classical KM, which depends on a single complete simple graph but may have natural frequencies depending on each node, without a rigorous mathematical guarantee very early in [7], and fully discussed for the case of equally placed natural frequencies with any odd node number $n\geq 3$ very recently in [34]. In particular, bifurcations and stability of synchronized solutions were proven to be very different between the classical KM and its CL. Moreover, bifurcations of completely synchronized solutions or twisted solutions in CLs for the KM of identical oscillators with two-mode interaction depending on two graphs or defined on nearest neighbor graphs with feedback control or not were discussed by the center manifold reduction technique [13], which is a standard one in dynamical systems theory [11, 32], in [35, 36, 37]. Similar CLs were also utilized for the KM with nonlocal coupling and a single or zero natural frequency in [10, 23, 25, 33].
The control problem of these coupled oscillator networks is important not only in theoretical interest but also in applications, and has attracted much attention [3, 6, 26]. For instance, coupled oscillator networks defined on multiple graphs were studied from the viewpoint of control [8, 9, 14, 35]: New graphs on which the coupled oscillators depend are added to avoid undesired synchronization since it may yield traffic congestion in networks and their collapse. On the other hand, feedback control of synchronized states in the KM on deterministic dense, random dense and random sparse graphs was studied numerically or theoretically in [30, 14, 37]. These results are also significant in applications since there often occurs such a situation as all coupled oscillators are desired to exhibit the same rhythmic motion like heart beats in contrast.
In this paper, we deal with the case of uniform graphs, i.e., $W(x,y)=p$, which may be deterministic dense, random dense or random sparse, and attempt to control the KM (1.1) so that each node exhibits the desired motion $V(t)=V_{1}t+V_{0}$, where $V_{1},V_{0}$ are constants, although more general desired motion can be treated similarly. So we add the control force $b_{1}\sin(V(t)-u_{i}^{n}(t))+b_{0}$ to each node as
$\frac{\mathrm{d}}{\mathrm{d}t}u_{i}^{n}(t)=$ $\omega_{i}^{n}+\frac{K}{n\alpha_{n}}\sum_{j=1}^{n}w_{ij}^{n}\sin\left(u_{j}^{n}(t)-u_{i}^{n}(t)\right)$
$+b_{1}\sin(V(t)-u_{i}^{n}(t))+b_{0},\quad i\in[n],$ (1.7)
where the constants $b_{1}$ and $b_{0}$ represent the feedback gain and stationary input, respectively. Here we include the stationary input $b_{0}$ for simplicity of analysis. The corresponding CL is given by
$\frac{\partial}{\partial t}u(t,x)=$ $\omega(x)+pK\int_{I}\sin(u(t,y)-u(t,x))\mathrm{d}y$
$+b_{1}\sin(V(t)-u(t,x))+b_{0}.$ (1.8)
A restricted case was studied and some numerical simulation results were provided in [14].
We see that the CL (1.8) has a particular solution
$u(t,x)=U(x)+V(t),\quad U(x)=\arcsin\left(\frac{\omega(x)-V_{1}+b_{0}}{pKC+b_{1}}\right),$ (1.9)
where the constant $C>0$ satisfies
$C=\int_{I}\sqrt{1-\left(\frac{\omega(x)-V_{1}+b_{0}}{pKC+b_{1}}\right)^{2}}\mathrm{d}x,$ (1.10)
if
$b_{0}=V_{1}-\int_{I}\omega(x)\mathrm{d}x$ (1.11)
and
$\sup_{x\in I}\,|\omega(x)-V_{1}+b_{0}|\leq pKC+b_{1}.$ (1.12)
In particular, the solution (1.9) is continuous in $t$ and $x$. Moreover, when the graphs $G_{n}$, $n\in\mathbb{N}$, are complete, i.e., $w_{ij}^{n}=p$, $i,j\in[n]$, the controlled Kuramoto model (CKM) (1.7) has a particular solution
$u_{i}^{n}(t)=U_{i}^{n}+V(t),\quad U_{i}^{n}=\arcsin\left(\frac{\omega_{i}^{n}-V_{1}+b_{0}}{pKC_{\mathrm{D}}+b_{1}}\right),$ (1.13)
where the constant $C_{\mathrm{D}}>0$ satisfies
$C_{\mathrm{D}}=\frac{1}{n}\sum_{i=1}^{n}\sqrt{1-\left(\frac{\omega_{i}^{n}-V_{1}+b_{0}}{pKC_{\mathrm{D}}+b_{1}}\right)^{2}}$ (1.14)
if
$b_{0}=V_{1}-\frac{1}{n}\sum_{i=1}^{n}\omega_{i}^{n}$ (1.15)
and
$\sup_{i\in[n]}\,|\omega_{i}^{n}-V_{1}+b_{0}|\leq pKC_{\mathrm{D}}+b_{1}.$ (1.16)
See Appendix A for the derivation of (1.9) and (1.13). In particular, as $b_{1}\to\infty$, the solutions (1.9) and (1.13), respectively, tend to
$u(t,x)=V(t)\quad\mbox{and}\quad u_{i}^{n}(t)=V(t),$
which coincide with the desired motions completely.
Specifically, we choose
$\omega(x)=a(x-\tfrac{1}{2}),$ (1.17)
where $a>0$ is a constant, as the frequency function as in [34], so that by (1.6)
$\omega_{i}^{n}=\frac{a}{2n}(2i-n-1),\quad i\in[n],$
i.e., the natural frequencies are uniformly spaced. In addition, by (1.11) and (1.15) $b_{0}=V_{1}$. We assume that the node number $n\geq 3$ and prove that for $a,pK>0$ fixed the solution (1.13) to the CKM (1.7) suffers a saddle-node bifurcation at some value of $b_{1}$ where a pair of stable and unstable ones born, and the solution (1.9) is always asymptotically stable in the CL (1.8) if it exists, as in the uncontrolled KM (1.1) and its CL (1.5) (see [34]). They are also proven to be the only synchronized solutions that are or can be stable in the CKM (1.7) and CL (1.8), respectively, when $b_{1}>0$ and $K>0$ is small enough for them not to exist for $b_{1},b_{0}=0$. Moreover, we show based on the results of [14, 34] that the solution (1.9) to the CL (1.8) behaves as it is an asymptotically stable one in the CKM (1.7) for $n>0$ sufficiently large, even if the graphs are random dense or sparse. In the companion paper [16], we discuss a similar control problem for the KM (1.1) with random natural frequencies.
The outline of this paper is as follows. In Section 2, we review the fundamental theoretical results of [14] including [34] for coupled nonlinear networks in the context of (1.7) and (1.8). In Section 3, we give our main results for synchronized solutions to the CKM (1.7) on complete graphs. The CKM (1.7) is transformed to an autonomous system of the same dimension, all equilibria there are found and their bifurcations and stability are detected or determined. In particular, it is proven that an equilibrium corresponding to the solution (1.13) in the CKM (1.7) suffers a saddle-node bifurcation where a pair of stable and unstable ones are born, and the other equilibria are unstable when $b_{1}$ is positive and $K>0$ is small enough for no equilibrium to exist for $b_{1},b_{0}=0$. In Section 4, we show that the solution (1.9) is always asymptotically stable if it exists and the other continuous and discontinuous synchronized solutions are unstable in the CL (1.8) when $b_{1}$ is positive and $K>0$ is small enough for no synchronized solution to exist for $b_{1},b_{0}=0$. In Section 6, we state the implication of the results of Section 4 in the CL (1.8) to the CKM (1.7) on random dense and sparse graphs, based on the fundamental theory in Section 2. Finally, we demonstrate the theoretical results by numerical simulations for the CKM (1.7) on complete simple, and uniform random dense and sparse graphs in Section 6.
2. Previous Fundamental Results
We review the fundamental theoretical results from [14, 34] in the context of (1.7) and (1.8). See Section 2 and Appendices A and B of [14] and Section 2 of [34] for more details including the proofs of the theorems stated below.
We begin with the initial value problem (IVP) of the CL (1.8). Let $g(x)\in L^{2}(I)$ and let $\mathbf{u}:\mathbb{R}\to L^{2}(I)$ stand for an $L^{2}(I)$-valued function on $\mathbb{R}$. We have the following from Theorem 2.1 of[14].
Theorem 2.1.
There exists a unique solution $\mathbf{u}(t)\in C^{1}(\mathbb{R};L^{2}(I))$ to the IVP of (1.8) with
$u(0,x)=g(x).$ (2.1)
Moreover, the solution depends continuously on $g$.
We next consider the IVP of the CKM (1.7) and discuss the convergence of solutions in (1.7) to those in the CL (1.8). Since the right-hand side of (1.7) is Lipschitz continuous in $u_{i}^{n}$, $i\in[n]$, we see by a fundamental result of ordinary differential equations (ODEs) (e.g., Theorem 2.1 of Chapter 1 of [4]) that the IVP of (1.7) has a unique solution. Given a solution $u_{n}(t)=(u_{1}^{n}(t),u_{2}^{n}(t),\ldots,u_{n}^{n}(t))$ to the IVP of the CKM (1.7), we define an $L^{2}(I)$-valued function $\mathbf{u}_{n}:\mathbb{R}\to L^{2}(I)$ as
$\mathbf{u}_{n}(t)=\sum_{i=1}^{n}u_{i}^{n}(t)\mathbf{1}_{I_{i}^{n}},$ (2.2)
[…]
Theorem 2.2.
If $\mathbf{u}_{n}(t)$ is the solution to the IVP of the CKM (1.7) with the initial condition
[…]
then for any $\tau>0$ we have
[…]
where $\mathbf{u}(t)$ represents the solution to the IVP of the CL (1.8) with (2.1).
We next discuss the stability of solutions to (1.7) and (1.8). We also obtain the following result, slightly modifying the proof of Theorem 2.5 in [14] (see also Theorem 2.3 of [34]).
Theorem 2.3.
Suppose that the CKM (1.7) and CL (1.8) have solutions $\bar{\mathbf{u}}_{n}(t)$ and $\bar{\mathbf{u}}(t)$, respectively, such that
[…]
for any $t\in[0,\infty)$. Then the following hold::\,::
  1. (i)
    If $\bar{\mathbf{u}}_{n}(t)$ is stable ((((resp. asymptotically stable)))) a.s. for $n>0$ sufficiently large, then $\bar{\mathbf{u}}(t)$ is also stable ((((resp. asymptotically stable);));) ;
  2. (ii)
    If $\bar{\mathbf{u}}(t)$ is stable, then for any $\varepsilon,\tau>0$, there exists $\delta>0$ such that for $n>0$ sufficiently large, if $\mathbf{u}_{n}(t)$ is any solution to the CKM (1.7) satisfying
[…]
    then
[…]
    Moreover, if $\bar{\mathbf{u}}(t)$ is asymptotically stable, then
[…]
    where $\mathbf{u}_{n}(t)$ is any solution to (1.7) such that $\mathbf{u}_{n}(0)$ is contained in the basin of attraction for $\bar{\mathbf{u}}(t)$.
Remark 2.4.
In the definition of stability and asymptotic stability of solutions to the CL (1.8), we cannot distinguish two solutions that are different only in a set with the Lebesgue measure zero. See also Section $2$ of [34].
Without assuming the existence of the solutions $\bar{\mathbf{u}}_{n}(t)$ to the CKM (1.7) satisfying (2.3), we have the following result as a corollary of Theorem 2.3 (see Corollary 2.6 of [34]).
Corollary 2.5.
If $\bar{\mathbf{u}}(t)$ is stable in the CL (1.8), then for any $\varepsilon,\tau>0$, there exists $\delta>0$ such that for $n>0$ sufficiently large, if $\mathbf{u}_{n}(t)$ is any solution to (1.7) satisfying
[…]
then
[…]
Moreover, if $\bar{\mathbf{u}}(t)$ is asymptotically stable, then
[…]
where $\mathbf{u}_{n}(t)$ is any solution to (1.7) such that $\mathbf{u}_{n}(0)$ is contained in the basin of attraction for $\bar{\mathbf{u}}(t)$.
Remark 2.6.
Corollary 2.5 implies that $\bar{\mathbf{u}}(t)$ behaves as if it is an (asymptotically) stable solution in the CKM (1.7), when $n>0$ is sufficiently large.
We have the following from Theorems 2.7 and 2.9 of [34].
Theorem 2.7.
Suppose that the hypothesis of Theorem 2.3 holds. Then the following hold::::
  1. (i)
    If $\mathbf{u}_{n}(t)$ is unstable a.s. for $n>0$ sufficiently large and no stable solution to the CKM (1.7) converges to $\mathbf{u}(t)$ a.s. as $n\to\infty$, then $\mathbf{u}(t)$ is unstable;
  2. (ii)
    If $\mathbf{u}(t)$ is unstable, then so is $\mathbf{u}_{n}(t)$ for $n>0$ sufficiently large.
Theorem 2.8.
If $\bar{\mathbf{u}}(t)$ is unstable, then for any $\varepsilon,\delta>0$ there exist $\tau,N>0$ such that for $n>N$
$\|\mathbf{u}_{n}(\tau)-\bar{\mathbf{u}}(\tau)\|>\varepsilon,$
where $\mathbf{u}_{n}(t)$ is a solution to the KM (1.7) satisfying
$\|\mathbf{u}_{n}(0)-\bar{\mathbf{u}}(0)\|<\delta.$
Remark 2.9.
  1. (i)
    Only under the hypothesis of Theorem 2.7, $\mathbf{u}(t)$ is not necessarily unstable even if $\mathbf{u}_{n}(t)$ is unstable for $n>0$ sufficiently large. Moreover, $\mathbf{u}(t)$ may be asymptotically stable even if $\mathbf{u}_{n}(t)$is unstable for $n>0$ sufficiently large. This behavior was proven to occur for the classical KM (1.1) on complete simple graphs and its CL (1.5) previously in [34]. We will see below that it also occurs for the CKM (1.7) on complete graphs and its CL (1.8).
  2. (ii)
    Theorem 2.8 implies that $\bar{\mathbf{u}}(t)$ behaves as if it is an unstable solution in the CKM (1.7), when $n>0$ is sufficiently large.
3. Controlled Kuramoto Model on Complete Graphs
We now take any integer $n\geq 3$ as the node number and consider the case in which the graph $G_{n}$ is deterministic and complete, i.e., $w_{ij}=p\in(0,1]$, $i,j\in[n]$, and $\alpha_{n}=1$. Let
$v_{i}=u_{i}^{n}-V(t),\quad i\in[n].$ (3.1)
Since we take $b_{0}=V_{1}$, Eq. (1.7) becomes
$\dot{v}_{i}=(2i-n-1)\nu+\frac{pK}{n}\sum_{j=1}^{n}\sin\left(v_{j}-v_{i}\right)-b_{1}\sin v_{i},\quad i\in[n],$ (3.2)
where $\nu=a/2n$. If Eq. (3.2) has an equilibrium at $v_{i}=\zeta_{i}$, $i\in[n]$, then
$u_{i}^{n}=\zeta_{i}+V(t),\quad i\in[n].$ (3.3)
is a solution to the CKM (1.7). Let $v=(v_{1},\ldots,v_{n})$ and $\zeta=(\zeta_{1},\ldots,\zeta_{n})$. If the equilibrium $v=\zeta$ is (asymptotically) stable, then so is the solution (3.3), and if a bifurcation of $v=\zeta$ occurs, then so does that of (3.3). Hence, we analyze the system (3.2) instead of the CKM (1.7) below. Our approaches are modifications of ones used in [34], in which the classical KM given by (1.1) with $p=1$ was analyzed when $n$ is any odd number with $n\geq 3$, although they are never straightforward.
3.1. Equilibria
Let $\sigma=\{\sigma_{i}\}_{i=1}^{n}$ be a sequence of length $n$ with $\sigma_{i}\in\{-1,1\}$, $i\in[n]$, and let
$\Sigma_{n}=\left\{\sigma=\{\sigma_{i}\}_{i=1}^{n}\mid\sigma_{i}\in\{-1,1\},i\in[n]\right\}.$
For each $\sigma\in\Sigma_{n}$, we define $C_{\mathrm{D}}^{\sigma}$ such that
$C_{\mathrm{D}}^{\sigma}=\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\sqrt{1-\left(\frac{(2i-n-1)\nu}{pKC_{\mathrm{D}}^{\sigma}+b_{1}}\right)^{2}},$ (3.4)
and write
$v_{i}^{\sigma}=\begin{cases}\phi_{i}&\mbox{if $\sigma_{i}=1$;}\\ \pi-\phi_{i}&\mbox{if $\sigma_{i}=-1$ and $\phi_{i}\geq 0$;}\\ -\phi_{i}-\pi&\mbox{if $\sigma_{i}=-1$ and $\phi_{i}<0$,}\end{cases}$ (3.5)
where
$\phi_{i}=\arcsin\left(\frac{(2i-n-1)\nu}{pKC_{\mathrm{D}}^{\sigma}+b_{1}}\right).$ (3.6)
As in Theorem 4.1 of [34], we can prove the following on the existence of equilibria in (3.2).
Theorem 3.1.
  1. (i)
    For each $\sigma\in\Sigma_{n}$, $v=v^{\sigma}\in\mathbb{T}^{n}:=\prod_{i=1}^{n}\mathbb{S}^{1}$ gives an equilibrium in (3.2), when $C_{\mathrm{D}}^{\sigma}$ satisfies (3.4). Moreover, no other equilibrium exists in (3.2) for $b_{1}\neq 0$.
  2. (ii)
    Fix $b_{1},a,pK>0$ and let $i\leq n/2$. If the equilibrium $v=v^{\sigma}$ with $\sigma_{i}=-1$ and $\sigma_{n-i+1}=1$ exists, then so does $v=v^{\bar{\sigma}}$ with $\bar{\sigma}_{i}=1$, $\bar{\sigma}_{n-i+1}=-1$ and $\bar{\sigma}_{j}=\sigma_{j}$, $j\in[n]\setminus\{i,n-i+1\}$, and vice versa.
Proof.
Let $\zeta=(\zeta_{1},\cdots,\zeta_{n})$ denote an equilibrium of (3.2). Recall that $b_{0}=V_{1}$. We begin with the following lemma.
Lemma 3.2.
We have
$\sum_{i=1}^{n}\sin\zeta_{i}=0.$
Proof.
Since the right-hand sides of (3.2) are zero when $v_{i}=\zeta_{i}$, $i\in[n]$, we sum up them to obtain the desired result. ∎
Let
$C_{\zeta}=\frac{1}{n}\sum_{i=1}^{n}\cos\zeta_{i}.$
Using Lemma 3.2 and (3.2), We have
$(2i-n-1)\nu-\left(pKC_{\zeta}+b_{1}\right)\sin\zeta_{i}=0.$ (3.7)
Hence, if $\phi_{i}\geq 0$ (resp. $\phi_{i}<0$), then $\zeta_{i}=\phi_{i}$ or $\pi-\phi_{i}$ (resp. $\zeta_{i}=\phi_{i}$ or $-\pi-\phi_{i}$), $i\in[n]$, where $\phi_{i}$ is given by (3.6) with $C_{\mathrm{D}}^{\sigma}=C_{\zeta}$. This yields part (i). By (3.4), we have $C_{\mathrm{D}}^{\sigma}=C_{\mathrm{D}}^{\bar{\sigma}}$, so that $\phi_{i}^{\sigma}=\phi_{i}^{\bar{\sigma}}$, $i\in[n]$. Hence, we obtain part (ii). ∎
For $\sigma\in\Sigma_{n}$, we define
$\chi^{\sigma}(\xi)=\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\sqrt{1-\left(\frac{2i-n-1}{n-1}\xi\right)^{2}},$
where $\xi\in[0,1]$. Letting $\xi=(n-1)\nu/|pKC_{D}^{\sigma}+b_{1}|$, we rewrite (3.4) as
$\frac{pK}{b_{1}}=\bar{\chi}^{\sigma}(\xi):=\frac{\xi}{\pm\beta-\xi\chi^{\sigma}(\xi)},\quad\beta=\frac{(n-1)\nu}{pK}>0,$ (3.8)
where the upper or lower is taken, depending on whether $pK\chi^{\sigma}(\xi)+b_{1}$ is positive or negative. From Theorem 3.1(i) we immediately obtain the following corollary.
Corollary 3.3.
Fix the value of $pK>0$. If $\xi\in(0,1]$ satisfies (3.8) for $\sigma\in\Sigma_{n}$ and $b_{1}>0$, then $v^{\sigma}$ given by (3.5) with
$\phi_{i}=\pm\arcsin\left(\frac{2i-n-1}{n-1}\xi\right),$
instead of (3.6), is an equilibrium in (3.2), where the upper or lower is taken, depending on whether $pK\chi^{\sigma}(\xi)+b_{1}$ is positive or negative.
Remark 3.4.
  1. (i)
    It is clear that $v^{\sigma}\neq v^{\hat{\sigma}}$ if $\sigma\neq\hat{\sigma}$, $\xi\neq 1$ and $b_{1}\neq 0$.
  2. (ii)
    From the result of [14, 34] we see that if $b_{1}=0$, then there exists a one-parameter family of equilibria in (3.2).
Figure 1. Function $\bar{\chi}^{\sigma}(\xi)$ for $n=4$: (a) and (b) $\beta=1$; (c) and (d) $0.1$. In plates (a) and (c) (resp. plates (b) and (d)) the sign ‘$+$’ (resp. ‘--’) is taken in (3.8). See the text for more details.
The graph of $\bar{\chi}^{\sigma}(\xi)$ is displayed for $n=4$ in Fig. 1. It is plotted as a solid line if $\bar{\chi}^{\sigma}(\xi)$ has a local maximum or minimum, and as a dotted or dashed line otherwise, depending on whether $\sigma_{1}\neq\sigma_{n}$ or not. The black and blue lines represent the graph for $(\sigma_{2},\sigma_{3})=(1,1)$ and $(-1,-1)$, respectively, while the red line for $(\sigma_{2},\sigma_{3})=(1,-1)$ or $(-1,1)$. We see that $\bar{\chi}^{\sigma}(\xi)$ is bounded if $\beta>0$ is sufficiently large and that it is unbounded if $\beta>0$ is sufficiently small. We prove the following.
Proposition 3.5.
Suppose that
$\beta=\frac{(n-1)\nu}{pK}>\max_{\xi\in[0,1]}\xi\chi^{\sigma}(\xi)\quad\mbox{with $\sigma_{i}=1$, $i\in[n]$}.$ (3.9)
Then the system (3.2) has no equilibrium when $b_{1}=0$.
Proof.
Assume that the hypothesis of the proposition holds. Then for any $\sigma\in\Sigma_{n}$
$\beta>\max_{\xi\in[0,1]}|\xi\chi^{\sigma}(\xi)|,$
so that $\bar{\chi}^{\sigma}(\xi)$ is bounded on $[0,1]$. This implies the desired result since the relation (3.8) does not hold near $b_{1}=0$. ∎
3.2. Bifurcations
We next discuss bifurcations of the equilibria in (3.2) detected by Theorem 3.1 and equivalently by Corollary 3.3. The relation (3.8) gives a branch of equilibria when $b_{1}$ is taken as a control parameter. As in Theorem 5.1 of [34], we prove the following on their bifurcations.
Theorem 3.6.
Fix the values of $a,pK>0$ and choose $b_{1}$ as a control parameter. Then the following hold.
  1. (i)
    The equilibrium $v^{\sigma}$ suffers a supercritical ((((resp. subcritical)))) saddle-node bifurcation at
    $b_{1}=\frac{pK}{\bar{\chi}^{\sigma}(\xi_{0})}$ (3.10)
    in (3.2) with $\xi=\xi_{0}$ on $(0,1)$ if and only if $\bar{\chi}^{\sigma}(\xi)$ has a local maximum ((((resp. minimum)))). In particular, if $|\bar{\chi}^{\sigma}(\xi)|$ is bounded on $[0,1]$ and
    $\sigma_{1},\sigma_{n}=1,\quad\bar{\chi}^{\sigma}(1)\geq 0\quad(\mbox{resp. }\sigma_{1},\sigma_{n}=-1,\quad\bar{\chi}^{\sigma}(1)\leq 0),$ (3.11)
    then a supercritical ((((resp. subcritical)))) saddle-node bifurcations occurs. Moreover, if $b_{1}>0$, $\sigma_{i}=1$, $i\in[n]$, and condition (3.9) holds, then $\bar{\chi}^{\sigma}(\xi)$ has a unique maximum and no local minimum, and $v^{\sigma}$ suffers only one saddle-node bifurcation.
  2. (ii)
    Let $v^{\sigma^{\pm\pm}}$ be four equilibria in (3.2) such that
    $\sigma_{i}^{++}=\sigma_{i}^{--}=\sigma_{i}^{+-}=\sigma_{i}^{-+},\quad i\neq 1,n,$
    and
    $\sigma_{1}^{++},\sigma_{1}^{+-}=1,\quad\sigma_{1}^{-+},\sigma_{1}^{--}=-1,\quad\sigma_{n}^{++},\sigma_{n}^{-+}=1,\quad\sigma_{n}^{+-}\sigma_{n}^{--}=-1.$
    If $\bar{\chi}^{\sigma}(1)\neq 0$, then a pitchfork bifurcation where $v^{\sigma^{++}}$ changes to $v^{\sigma^{--}}$ and where $v^{\sigma^{+-}}$ and $v^{\sigma^{-+}}$ are born occurs at
    $b_{1}=\frac{pK}{\bar{\chi}^{\sigma}(1)},$ (3.12)
    where any of $\sigma^{\pm\pm}$ may be chosen as $\sigma$. Moreover, the bifurcation is super- or subcritical if
    $\frac{\mathrm{d}\bar{\chi}^{\sigma}}{\mathrm{d}\xi}(1)\quad\mbox{with $\sigma=\sigma^{+-}$ and $\sigma^{-+}$}$ (3.13)
    is positive or negative, where any of $\sigma^{+-}$ and $\sigma^{-+}$ may be chosen as $\sigma$.
Proof.
We first easily see via Corollary 3.3 that if $\xi\in(0,1)$ satisfies (3.8). Hence, if $\bar{\chi}^{\sigma}(\xi)$ has a local maximum (resp. minimum) at $\xi=\xi_{0}$ on $(0,1)$, then no value of $\xi$ satisfies (3.8) for values of $b_{1}$ that are smaller (resp. larger) than and close to (3.10) but two values of $\xi$ satisfy it for values of $b_{1}$ that are larger (resp. smaller) than and close to (3.10), so that the equilibrium $v^{\sigma}$ suffers a supercritical ((((resp. subcritical)))) saddle-node bifurcation there. If $\bar{\chi}^{\sigma}(\xi)$ is bounded on $[0,1]$ and condition (3.11) (resp. (3.12)) holds, then $\bar{\chi}^{\sigma}(\xi)$ has a local maximum (resp. minimum), since $\bar{\chi}^{\sigma}(0)=0$ and
$\frac{\mathrm{d}\bar{\chi}^{\sigma}}{\mathrm{d}\xi}(\xi)=\frac{\displaystyle\pm\beta+\xi^{2}\frac{\mathrm{d}\chi^{\sigma}}{\mathrm{d}\xi}(\xi)}{(\pm\beta-\xi\chi^{\sigma}(\xi))^{2}}\to-\infty\quad(\mbox{resp. }+\infty)$. + ∞ ) (3.14)
as $\xi\to 1$. Note that
$\frac{\mathrm{d}\chi^{\sigma}}{\mathrm{d}\xi}(\xi)$
$=-\frac{1}{n}\sum_{i=1}^{n}\sigma_{i}\left(\frac{2i-n-1}{n-1}\right)^{2}\xi\Bigg{/}\sqrt{1-\left(\frac{2i-n-1}{n-1}\xi\right)^{2}}\to-\infty\quad(\mbox{resp. }+\infty)$. + ∞ )
as $\xi\to 1$ under the conditions.
Assume that $b_{1}>0$, $\sigma_{i}=1$, $i\in[n]$, and condition (3.9) holds. Since $pK\chi^{\sigma}(\xi)+b_{1}>0$, we have
$\bar{\chi}^{\sigma}(\xi)=\frac{\xi}{(\beta-\xi\chi^{\sigma}(\xi))^{2}}>0$
and
$\frac{\mathrm{d}\bar{\chi}^{\sigma}}{\mathrm{d}\xi}(\xi)=\frac{\displaystyle\beta-\frac{\xi}{n}\sum_{i=1}^{n}\left(\frac{2i-n-1}{n-1}\xi\right)^{2}\Bigg{/}\sqrt{1-\left(\frac{2i-n-1}{n-1}\xi\right)^{2}}}{(\beta-\xi\chi^{\sigma}(\xi))^{2}}.$ (3.15)
The numerator in (3.15) is monotonically decreasing on $(0,1)$ and positive at $\xi=0$ and goes to $\infty$- ∞ as $\xi\to 1$ while the denominator is always positive. Hence, $\bar{\chi}^{\sigma}(\xi)$ has a unique local maximum at which its derivative has a unique zero, and it has no local minimum. Thus, we obtain part (i).
On the other hand, assume that $\bar{\chi}^{\sigma}(1)\neq 0$. Then at $\xi=1$, the four equilibria $v^{\sigma^{\pm\pm}}$ coincide and the corresponding functions $\bar{\chi}^{\sigma^{\pm\pm}}(\xi)$ have the same value. We see that $\sigma^{++}$ and $\sigma^{--}$, respectively, satisfy the first and second equations of (3.14) as $\xi\to 1$, and that $v^{\sigma^{+-}}$ and $v^{\sigma^{-+}}$ exist in a pair by Theorem 3.1(ii) for values of $b_{1}$ that are larger or smaller than and close to (3.12), depending on whether the quantity (3.13), which takes the same value for both of $\sigma=\sigma^{+-}$ and $\sigma^{-+}$, is positive or negative. So we obtain part (ii). ∎
As in Proposition 4.4 of [34], we also have the following.
Proposition 3.7.
For any $\sigma\in\Sigma_{n}$, the equilibrium $v^{\sigma}$ suffers no Hopf bifurcation.
Proof.
We compute each element of the Jacobian matrix $A$ for the vector field of (3.2) as
$A_{ij}=\begin{cases}-\frac{pK}{n}\sum_{j=1,j\neq i}^{n}\cos\left(v_{j}-v_{i}\right)-b_{1}\cos v_{i}&\mbox{if $i=j$};\\ \frac{pK}{n}\cos\left(v_{j}-v_{i}\right)&\mbox{if $i\neq j$}.\end{cases}$ (3.16)
Thus, the matrix $A$ is a symmetric and consequently has only real eigenvalues. This implies the desired result. ∎
Remark 3.8.
From Remark 3.4(i) and Proposition 3.7 we see that no other bifurcation than ones detected in Theorem 3.6 occurs for equilibria in the system (3.2) if $b_{1}\neq 0$.
3.3. Stability
We now discuss the stability of the equilibria detected by Theorem 3.1 and equivalently by Corollary 3.3 in (3.2). As in Theorem 6.1 of [34], we prove the following on their stability.
Theorem 3.9.
Suppose that $b_{1}>0$ and consition (3.9) holds. Then the following hold::::
  1. (i)
    The equilibrium $v^{\sigma}$ with $\sigma_{i}=1$ for all $i\in[n]$ is asymptotically stable if $\xi=(n-1)\nu/|pKC_{\mathrm{D}}^{\sigma}+b_{1}|<\xi_{0}$ and unstable if $\xi>\xi_{0}$, where $\xi_{0}$ is the unique local maximum of $\bar{\chi}^{\sigma}(\xi)$ detected in Theorem 3.6(i).
  2. (ii)
    The equilibrium $v^{\sigma}$ is unstable if $\sigma_{i}=-1$ for some $i\in[n]$.
Proof.
Suppose that the hypotheses of the theorem hold. Fix $\sigma\in\Sigma_{n}$ and increase $\xi\in(0,1)$ along a branch of equilibria, $v^{\sigma}$ with $\bar{\chi}^{\sigma}(\xi)$. We write $v^{\sigma}(\xi)=v^{\sigma}\in\mathbb{T}^{n}$ and $f(v;b_{1})$ for the vector field of (3.2). By Proposition 3.5 the branch $b_{1}=pK/\bar{\chi}^{\sigma}(\xi)$ does not intersect $b_{1}=0$. We begin with the following lemma.
Lemma 3.10.
Suppose that the multiplicity of the zero eigenvalue of the Jacobian matrix $\mathrm{D}_{v}f(v^{\sigma}(\xi);b_{1})$ with $\bar{\chi}^{\sigma}(\xi)$, $\sigma\in\Sigma_{n}$, changes at $\xi=\xi_{*}$ when $\xi$ increases on $(0,1)$. Then $\bar{\chi}^{\sigma}(\xi)$ has an extremum at $\xi=\xi_{*}$ and the change of the multiplicity is one at most.
Proof.
We apply arguments similar to those used in the proof of Lemma $6.2$ of [34].
Since $\mathrm{D}f(v^{\sigma}(\xi);b_{1})$ is symmetric, its eigenvalues are real and geometrically simple, as shown in the proof of Proposition 3.7. In particular, if the eigenvalues change their signs, then they must become zero.
Assume that $\mathrm{D}f(v^{\sigma}(\xi);b_{1})$ has a simple zero eigenvalue when Eq. (3.8) holds with $\xi=\xi_{*}\in(0,1)$. Let $\bar{e}\in\mathbb{R}^{n}$ denote the associated eigenvector. Then the equilibrium $v^{\sigma}(\xi_{*})$ has a one-dimensional center manifold [11, 19, 32] on which the system (3.2) reduces to
$\dot{v}_{\mathrm{c}}=c_{1}v_{\mathrm{c}}^{j}+\cdots+c_{2}(b_{1}-b_{1*})+\cdots,\quad v_{\mathrm{c}}\in\mathbb{R},$ (3.17)
where $b_{1*}=K/\bar{\chi}^{\sigma}(\xi_{*})$, $c_{1},c_{2}\in\mathbb{R}$ are constants and $j>1$ is an even integer, since the eigenvalue does not change its sign if $j$ is odd. If $\bar{e}$ is linearly independent of $(\mathrm{d}v^{\sigma}/\mathrm{d}\xi)(\xi_{*})$, then there exists a different equilibrium from $v^{\sigma}(\xi)$ close to $\xi=\xi_{*}$, but coincide with it at $\xi=\xi_{*}$. This contradicts that the equilibrium $v^{\sigma}(\xi_{*})$ is isolated for $\xi\in(0,1)$ by Remark 3.4(i). Hence, we can take $\bar{e}=(\mathrm{d}v^{\sigma}/\mathrm{d}\xi)(\xi_{*})$, so that by (3.8)
$\frac{\mathrm{d}\bar{\chi}^{\sigma}}{\mathrm{d}\xi}(\xi_{*})=0$
since
$~{}\frac{\mathrm{d}}{\mathrm{d}\xi}f\left(v^{\sigma}(\xi);\frac{pK}{\bar{\chi}^{\sigma}(\xi)}\right)\bigg{|}_{\xi=\xi_{*}}$
$=\mathrm{D}_{v}f(v^{\sigma}(\xi);b_{1*})\frac{\mathrm{d}v^{\sigma}}{\mathrm{d}\xi}(\xi_{*})-\frac{\partial f}{\partial b_{1}}(v^{\sigma}(\xi);b_{1*})\frac{pK}{\left(\bar{\chi}^{\sigma}(\xi)\right)^{2}}\frac{\mathrm{d}\bar{\chi}^{\sigma}}{\mathrm{d}\xi}(\xi)=0$
Similarly, we can show that $\mathrm{D}f(v^{\sigma}(\xi);b_{1})$ never has a non-simple zero eigenvalue even when $(\mathrm{d}\bar{\chi}^{\sigma}/\mathrm{d}\xi)=0$, since the equilibrium $v^{\sigma}(\xi)$ is isolated for $\xi\in(0,1)$. Thus, we obtain the desired result. ∎
We turn to the proof of Theorem 3.9. Taking the limit $\xi=(n-1)\nu/|pKC_{\mathrm{D}}^{\sigma}+b_{1}|\to 0$, we have $b_{1}\to+\infty$ and $\phi_{i}\to 0$ in (3.6), so that by (1.13)
$v_{i}^{\sigma}\to\begin{cases}0&\mbox{if $\sigma_{i}=1$;}\\ \pi\mbox{ or }-\pi&\mbox{if $\sigma_{i}=-1$,}\end{cases}$
since the branch $b_{1}=pK/\bar{\chi}^{\sigma}(\xi)$ does not intersect $b_{1}=0$. Let $n_{+}$ and $n_{-}$ be, respectively, the numbers of $\sigma_{i}=1$ and $-1$, $i\in[n]$. The Jacobian matrix $A=\mathrm{D}f(v^{\sigma}(\xi);b_{1})$ is written as
$A=b_{1}A_{0}+O(1),\quad b_{1}\to\infty,$
as $\xi\to 0$, where $A_{0}$ is an $n\times n$ diagonal matrix whose $n_{-}$ and $n_{+}$ diagonal elements are $1$ and $-1$, respectively. We immediately obtain the following.
Lemma 3.11.
Let $b_{1}>0$ be sufficiently large. Then the numbers of positive and negative eigenvalues of the matrix $A$ are $n_{-}$ and $n_{+}$, respectively. In particular, the matrix $A$ has no positive eigenvalue if and only if $n_{-}=n$, i.e., $n_{+}=0$.
Let $\sigma_{i}=1$, $i\in[n]$, i.e., $n_{-}=0$. Since by Theorem 3.6(i) $\bar{\chi}^{\sigma}(\xi)$ has an extremum only at $\xi=\xi_{0}$, it follows from Lemma 3.10 that only one eigenvalue of $\mathrm{D}f(v;b_{1})$ with $b_{1}=pK/\bar{\chi}^{\sigma}(\xi)$ changes its sign once and at $\xi=\xi_{0}$ when $\xi$ increases from zero to one. Noting that all of its eigenvalues are negative near $\xi=0$ by Lemma 3.11, we obtain part (i) .
We turn to the proof of part (ii). We compute
$\frac{\mathrm{d}v_{i}^{\sigma}}{\mathrm{d}\xi}(\xi)=\pm\sigma_{i}\left(\frac{2i-n-1}{n-1}\right)\bigg{/}\sqrt{1-\left(\frac{2i-n-1}{n-1}\xi\right)^{2}}$
and
$\frac{\partial f_{i}}{\partial b_{1}}(v^{\sigma}(\xi);b_{1})=\mp\sigma_{i}\left(\frac{2i-n-1}{n-1}\right)$
for $i\in[n]$ when $b_{1}=pK/\bar{\chi}^{\sigma}(\xi)$, where $f_{i}(v;b_{1})$ is the $i$th element of $f(v;b_{1})$ and the upper or lower sign is taken simultaneously, depending on whether $pK\chi^{\sigma}(\xi)+b_{1}$ is positive or not.
Figure 2. Bifurcation diagrams of equilibria in (3.17): (a) $c_{1}/c_{2}<0$; (b) $c_{1}/c_{2}>0$.
Suppose that $(\mathrm{d}\bar{\chi}^{\sigma}/\mathrm{d}\xi)(\xi)$ has a zero at $\xi=\xi_{*}$ and let $b_{1*}=aK/\bar{\chi}^{\sigma}(\xi_{*})>0$ be bounded. Then the Jacobian matrix $\mathrm{D}_{v}f(v^{\sigma}(\xi_{*});b_{1*})$ has a simple zero eigenvalue and $(\mathrm{d}v/\mathrm{d}\xi)(\xi_{*})$ is the associated eigenvector. The equilibrium $v^{\sigma}(\xi_{*})$ has a one-dimensional center manifold on which the system (3.2) reduces to (3.17). See the proof of Lemma 3.10. In addition, $c_{1}/c_{2}$ is negative or positive, whether $\bar{\chi}^{\sigma}(\xi)$ has a local maximum or minimum at $\xi=\xi_{*}$, since by Theorem 3.6(i) a supercritical or subcritical saddle-node bifurcation occurs. See Fig. 2. Moreover, substituting $v=(\mathrm{d}v^{\sigma}/\mathrm{d}\xi)(\xi_{*})v_{\mathrm{c}}+v^{\sigma}(\xi_{*})$ into (3.2) and taking the inner product of the resulting equation with $(\mathrm{d}v/\mathrm{d}\xi)(\xi_{*})$, we have
$c_{2}=\frac{\mathrm{d}v_{i}^{\sigma}}{\mathrm{d}\xi}(\xi_{*})\cdot\frac{\partial f_{i}}{\partial b_{1}}(v^{\sigma}(\xi_{*});b_{1*})\bigg{/}\biggl{|}\frac{\mathrm{d}v_{i}^{\sigma}}{\mathrm{d}\xi}(\xi_{*})\biggr{|}^{2}<0$
where the dot $\cdot$”represents the standard inner product, since
$\frac{\mathrm{d}v_{i}^{\sigma}}{\mathrm{d}\xi}(\xi_{*})\cdot\frac{\partial f_{i}}{\partial b_{1}}(v^{\sigma}(\xi_{*});b_{1*})=$ $-\sum_{i=1}^{n_{0}}\left(\frac{i-n_{0}-1}{n_{0}}\right)^{2}\bigg{/}\sqrt{1-\left(\frac{i-n_{0}-1}{n_{0}}\xi\right)^{2}}.$
We have
$v_{\mathrm{c}}=\frac{\mathrm{d}v^{\sigma}}{\mathrm{d}\xi}(\xi_{*})\cdot(v-v^{\sigma}(\xi_{*}))\bigg{/}\left|\frac{\mathrm{d}v^{\sigma}}{\mathrm{d}\xi}(\xi_{*})\right|^{2},$
so that $v_{c}$ is negative or positive for $v=v^{\sigma}(\xi)$ near $\xi=\xi_{*}$, depending on whether $\xi$ is less or larger than $\xi_{*}$. Hence, we see via (3.17) that a negative (resp. positive) eigenvalue of $\mathrm{D}_{v}f(v^{\sigma}(\xi);b_{1})$ with $b_{1}=pK/\bar{\chi}^{\sigma}(\xi)$ becomes positive (resp. negative) when $\bar{\chi}^{\sigma}(\xi)$ has a local maximum (resp. minimum) at $\xi=\xi_{*}$. Thus, we obtain the desired result. ∎
Remark 3.12.
From the proof of Theorem 3.9 we see that the equilibrium $v^{\sigma}$ with $\sigma_{i}=-1$, $i\in[n]$, is stable for $b_{1}<0$ with $|b_{1}|\gg 1$.
By the relation (3.1) we have
$u_{i}^{n}=v_{i}^{\sigma}+V(t)$ (3.18)
as a synchronized solution in the CKM (1.7) for each $\sigma\in\Sigma_{n}$. In particular, when $\sigma_{i}=1$, $i\in[n]$, we have
$u_{i}^{n}=\arcsin\left(\frac{2i-n-1}{n}\xi\right)+V(t),\quad\xi=\frac{n\nu}{pKC_{\mathrm{D}}^{\sigma}+b_{1}},$
which coincides with (1.13). The synchronized solution (3.18) suffers such bifurcations as detected in Theorem 3.6 and has the same stability type as the equilibrium $v^{\sigma}$ in (3.2), which is determined in Theorem 3.9. In particular, it is stable for $\xi<\xi_{*}$ and unstable for $\xi>\xi_{*}$ if $\sigma_{i}=1$ for $i\in[n]$, and always unstable if $\sigma_{i}=-1$ for some $i\in[n]$, under the hypotheses of Theorem 3.9.
4. Continuum Limits
We turn to the CL (1.8) with the frequency function (1.17). The synchronized continuous solution (1.9) becomes
$u(t,x)=U(x)+V(t),\quad U(x)=\arcsin\left(\frac{a(x-\tfrac{1}{2})}{KC}+b_{1}\right),$ (4.1)
where the constant $C>0$ satisfies
$C=$ $\frac{pKC+b_{1}}{a}\Biggl{(}\arcsin\left(\frac{a}{2(pKC+b_{1})}\right)$
$+\frac{a}{2(pKC+b_{1})}\sqrt{1-\left(\frac{a}{2(pKC+b_{1})}\right)^{2}}\Biggr{)},$ (4.2)
which follows from (1.10). Using the arguments in Appendix A (see Eq. (A.2) especially), we easily see that the CL (1.8) has another continuous solution
$u(t,x)=\pi-U(x)+V(t),$ (4.3)
where the constant $C$ also satisfies (4.2).
Proposition 4.1.
Let $pK/a<2/\pi$. The two continuous solutions (4.1) and (4.3) exist in the CL (1.8) if
$b_{1}\geq\tfrac{1}{2}a-\tfrac{1}{4}\pi pK.$ (4.4)
Proof.
Letting
$\varphi(\eta)=\frac{\beta_{0}-\left(\arcsin\eta+\eta\sqrt{1-\eta^{2}}\right)}{\eta},\quad\beta_{0}=\frac{a}{pK},$
we rewrite (4.2) as
$b_{1}=\tfrac{1}{2}pK\varphi\left(\frac{a}{2(pKC+b_{1})}\right).$
Since $\varphi(\eta)$ is monotonically decreasing on $(0,1)$ and
$\lim_{\eta\to 0}\varphi(\eta)=+\infty,\quad\varphi(1)=\beta_{0}-\tfrac{1}{2}\pi,$
we see that Eq. (4.2) has a solution $C>0$ if
$b_{1}\geq\tfrac{1}{2}pK\left(\beta_{0}-\tfrac{1}{2}\pi\right),$
which yields (4.4). This implies the desired result. ∎
Remark 4.2.
As shown in Section $7$ of [34], if $pK/a<2/\pi$, then such continuous synchronized solutions as (4.1) and (4.3) does not exist in the CL (1.8) with $b_{1},b_{0}=0$, i.e., in the CL (1.5).
The CL (1.8) also has infinitely many discontinuous solutions of the form
$u(t,x)=\begin{cases}U(x)+V(t)&\mbox{for $x\in[0,1]\setminus\hat{I}$;}\\ \pi-U(x)+V(t)&\mbox{for $x\in\hat{I}_{j}^{+}$, $j\in[m_{+}]$;}\\ -U(x)-\pi+V(t)&\mbox{for $x\in\hat{I}_{j}^{-}$, $j\in[m_{-}]$,}\end{cases}$ (4.5)
where $m_{\pm}$ are nonnegative integers and may be infinite, $\hat{I}_{j}^{\pm}\subset[0,1]$, $j\in[m_{\pm}]$, are intervals such that $\hat{I}_{j}^{-}\in[0,\tfrac{1}{2}]$, $\hat{I}_{j}^{+}\in[\tfrac{1}{2},1]$ and $\hat{I}_{j}^{-}\cap\hat{I}_{k}^{-},\hat{I}_{j}^{+}\cap\hat{I}_{k}^{+}=\emptyset$,
$\hat{I}=\bigcup_{j=1}^{m_{-}}\hat{I}_{j}^{-}\cup\bigcup_{j=1}^{m_{+}}\hat{I}_{j}^{+}\neq\emptyset,$
and the constant $C$ in $U(x)$ satisfies
$C=\int_{[0,1]\setminus\hat{I}}\sqrt{1-\left(\frac{a(x-\tfrac{1}{2})}{pKC+b_{1}}\right)^{2}}\mathrm{d}x-\int_{\hat{I}}\sqrt{1-\left(\frac{a(x-\tfrac{1}{2})}{pKC+b_{1}}\right)^{2}}\mathrm{d}x.$
Here the interiors of $\hat{I}_{j}^{\pm}$, $j\in[m_{\pm}]$, may be empty. We see that the one-parameter family of synchronized solutions given by (3.3) converge to (4.5) when
$\sigma_{i}=\begin{cases}1&\mbox{if $i/n\in[0,1]\setminus\hat{I}$;}\\ -1&\mbox{if $i/n\in\hat{I}$.}\end{cases}$ (4.6)
Using Theorems 2.3 and 2.7, we prove the following
Theorem 4.3.
Let $pK/a<2/\pi$ and $b_{1}>0$. Then the following hold::::
  1. (i)
    The continuous solution (4.1) is asymptotically stable while the continuous (4.3) is unstable when condition (4.4) holds;
  2. (ii)
    The discontinuous solution (4.5) with $\hat{I}\neq\emptyset$ is unstable if it exists and is different from (4.1) in the sense of $L^{2}(I)$.
Proof.
The synchronized solution (3.18) with $\sigma_{i}=1$ (resp. $\sigma_{i}=-1$), $i\in[n]$, which is asymptotically stable for $\xi<\xi_{0}$ (resp. unstable for $\xi\in(0,1)$) in the CKM (1.7) by Theorem 3.9(i), converges to the continuous solution (4.1) (resp. (4.3)) as $n\to\infty$. Recall that $\bar{\chi}^{\sigma}(\xi)$ has a unique extremum (local maximum) at $\xi=\xi_{0}$ (see Theorem 3.6(i)).
Let $\sigma_{i}=1$, $i\in[n]$. Since
$\chi^{\sigma}(\xi)\to 2\int_{0}^{1}\sqrt{1-\xi^{2}x^{2}}\,\mathrm{d}x=\frac{\xi\sqrt{1-\xi^{2}}+\arcsin\xi}{2\xi},$
we have
$\bar{\chi}^{\sigma}(\xi)=\frac{\xi}{\beta-\xi\chi^{\sigma}(\xi)}\to\frac{2\xi}{2\beta-(\xi\sqrt{1-\xi^{2}}+\arcsin\xi)}=:\bar{\chi}_{0}(\xi).$
Noting that
$\frac{\mathrm{d}}{\mathrm{d}\xi}\left(\xi\sqrt{1-\xi^{2}}+\arcsin\xi\right)=2\sqrt{1-\xi^{2}}>0,$
we see that $\bar{\chi}_{0}(\xi)$ is monotonically increasing on $(0,1)$. This implies that $\xi_{0}\to 1$ as $n\to\infty$. Using Theorem 2.3(i), we obtain part (i).
We turn to the proof of part (ii). The synchronized solution (3.18) with (4.6), which is unstable by Theorem 3.9(ii), converges to the discontinuous solution (4.1) in $L^{2}(I)$ as $n\to\infty$. Hence, by Theorem 2.7, if it is different from (4.1) in the sense of $L^{2}(I)$, then the discontinuous solution (4.5) with $\hat{I}\neq\emptyset$ is unstable. ∎
Remark 4.4.
[…]
From Theorems 3.6 and 3.9 we obtain the following results for the CKM (1.7) on complete graphs: The synchronized solution (3.3) for $v^{\sigma}$ with $\sigma_{i}=1$, $i\in[n]$, suffers a saddle-node bifurcation at $\xi=\xi_{0}$, i.e., $b_{1}=pK/\bar{\chi}^{\sigma}(\xi_{0})$, and it turns unstable from stable there. In addition, it suffers a pitchfork bifurcation at $\xi=1$, i.e., $b_{1}=pK/\bar{\chi}^{\sigma}(1)$, and it changes to a synchronized solution with $\sigma_{1}=\sigma_{n}=-1$ and $\sigma_{j}=1$, $i\in[n-1]\setminus\{1\}$, while the two synchronized solutions with $\sigma_{1}=-1$, $\sigma_{n}=1$ or $\sigma_{1}=1$, $\sigma_{n}=-1$ and $\sigma_{j}=1$, $i\in[n-1]\setminus\{1\}$, are born. When $n\to\infty$, the saddle-node and pitchfork bifurcations collide, since $\xi_{0}\to 1$ as shown in the proof of Theorem 4.3. Moreover, these four solutions converge to four solutions that are regarded as the same one in the sense of $L^{2}(I)$ and asymptotically stable as an $L^{2}(I)$ solution in the CL (1.8) by Theorem 4.3. More generally, if $\sigma_{i}=1$ except for a fixed finite positive number of $i\in[n]$, then by Theorem 4.3(ii) the synchronized solution (3.3) for $v^{\sigma}$ is unstable but converges to the asymptotically stable solution (4.1) in $L^{2}(I)$ as $n\to\infty$. Thus, bifurcation behavior in the CL (1.8) is very subtle, compared with such finite-dimensional dynamical systems as the CKM (1.7). Such behavior was previously detected for the KM (1.1) on complete simple graphs in [34].
5. Controlled Kuramoto Model on Uniform Random Graphs
We turn to the CKM (1.7) on uniform random dense and sparse graphs given by (1.2) or (1.3), respectively.
The second term in the right-hand side of (1.7) is very complicated, compared with the case of complete graphs, although the coefficient $w_{ij}$ is one or zero for all $i,j\in[n]$. So it is very difficult even to obtain equilibria analytically. However, the relationships of the CL (1.8) with the CKM (1.7) stated in Section 2 allow us to describe fundamental dynamical behavior in the CKM (1.7), based on the results of Section 4 for the CL (1.8).
Let $\bar{\mathbf{u}}(t)$ (resp. $\hat{\mathbf{u}}(t)$) denote solutions (4.1) (resp. (4.3) or (4.5)) to the CL (1.8), such that $\bar{\mathbf{u}}(t)$ and $\hat{\mathbf{u}}(t)$ are different in $L^{2}(I)$. Using Corollary 2.5 and Theorems 2.8 and 4.3, we prove the following.
Theorem 5.1.
  1. (i)
    For any $\varepsilon,\tau>0$ there exists $\delta,N>0$ such that for any $n>N$ if $\mathbf{u}_{n}(t)$ is any solution to the CKM (1.7) satisfying
[…]
    then
[…]
    Moreover,
[…]
  2. (ii)
    Let For any $\varepsilon,\delta>0$ there exist $\tau,N>0$ such that for $n>N$
[…]
    where $\mathbf{u}_{n}(t)$ denote a solution to the KM (1.7) satisfying
    $\|\mathbf{u}_{n}(0)-\bar{\mathbf{u}}(0)\|<\delta.$
Proof.
By Theorem 4.3, $\bar{\mathbf{u}}(t)$ and $\hat{\mathbf{u}}(t)$ are, respectively, asymptotically stable and unstable solutions in the CL (1.8). Hence, parts (i) and (ii) immediately follow from Corollary 2.5 and Theorem 2.8, respectively. ∎
Thus, the solution (4.1) behaves (resp. the solutions (4.3) and (4.5) behave) as if it is an asymptotically stable one (resp. they are unstable ones) in the CKM (1.7) on uniform random dense and sparse graphs. Similar behavior was previously numerically observed or theoretically detected for the KMs defined on some random dense and sparse graphs and their CLs including the KM (1.1) and CL (1.5) when the natural frequencies are deterministic or random, in [14, 35, 36, 37, 38].
6. Numerical Simulations
Figure 3. Pixel pictures of sampled weighted matrices for the random undirected graphs given by $w_{ij}^{n}=1$, $i,j\in[n]$ with probability (6.1) and (6.2) for $n=1000$: (b) Dense graph with $p=0.5$; (c) Sparse graph with $p=0.5$ and $\gamma=0.3$. The color of the corresponding pixel is blue if $w_{ij}=1$ and it is light blue otherwise.
We finally present numerical simulation results for the CKM (1.7). We treat the following three cases for the graph $G_{n}$:
  • (i)
    Complete simple graph, which is deterministic dense with $p=1$;
  • (ii)
    Random undirected dense graph in which $w_{ij}^{n}=1$ with probability
    $\mathbb{P}(i\sim j)=p,\quad i,j\in[n];$ (6.1)
  • (iii)
    Random undirected sparse graph in which $w_{ij}^{n}=1$ with probability
    $\mathbb{P}(i\sim j)=n^{-\gamma}p,\quad i,j\in[n],$ (6.2)
    where $\gamma\in(0,\tfrac{1}{2})$. Recall that $\alpha_{n}=n^{-\gamma}$ and note that $\alpha_{n}^{-1}=n^{\gamma}>1$ for any $n>1$.
We specifically take $p=0.5$ in case (ii) and $p=0.5$ and $\gamma=0.3$ in case (iii). Figures 3(a) and (b) provide numerically computed samples of the weight matrices for the random undirected dense and sparse graphs in cases (ii) and (iii), respectively, with $n=1000$, $p=0.5$ and $\gamma=0.3$.
We carried out numerical simulations for the CKM (1.7) on the three types of graphs with $n=1000$, $V_{1}=b_{0}=1$ and $V_{0}=1$ using the DOP853 solver [12]. We also chose $K=0.5$ and $a=1$ for case (i), and $K=0.5$ and $a=0.5$ for cases (ii) and (iii). The continuous solution (4.1) does not exist in the CL (1.8) with $b_{1},b_{0}=0$, in cases (i)-(iii) since $pK/a=0.5<2/\pi=0.63661\ldots$ (see Remark 4.2). The initial values $u_{i}^{n}(0)$, $i\in[n]$ were independently randomly chosen according to the uniform distribution on $[-\pi,\pi]$.
Figure 4. Numerical simulation results of the CKM (1.7) with $n=1000$, $K=0.5$, $V_{1},b_{0}=1$, $V_{0}=1$ and $b_{1}=0.2$: (a) $(a,p)=(1,1)$ in case (i); (b) $(0.5,0.5)$ in case (ii); (c) $(a,p,\gamma)=(0.5,0.5,0.3)$ in case (iii). The time-history of every 100th node (from 50th to 950th) is plotted with different colors.
Figures 4(a), (b) and (c) show the time-histories of every 100th node (from 50th to 950th) in cases (i), (ii) and (iii), respectively, for $b_{1}=0.2$, in which the asymptotically stable solution (4.1) exists in the CL (1.8). The ordinates represent the deviations of the responses from the desired motion, $u_{i}^{n}(t)-V(t)$. We observe that the responses rapidly converge to the synchronized solution around the desired motion, as predicted theoretically in Sections 3 and 5.
Figure 5. Deviations of steady state responses from the desired motion in the CKM (1.7) with $n=1000$, $K=0.5$, $V_{1},b_{0}=1$, $V_{0}=1$ and $b_{1}=0.2$: (a) $(a,p)=(1,1)$ in case (i); (b) $(0.5,0.5)$ in case (ii); (c) $(a,p,\gamma)=(0.5,0.5,0.3)$ in case (iii). Here $u_{i}^{n}(t)-V(t)$, $i\in[n]$, with $t=100$ are plotted as red dots. The blue line represents the corresponding theoretical predictions computed from the synchronized solution (4.1) in the CL (1.8).
In Figs. 5(a), (b) and (c), the deviations of the responses from the desired motion at time $t=100$, which may be regarded as the steady states from Fig. 4, are plotted for cases (i), (ii) and (iii), respectively. Here the same initial condition and value of $b_{1}$ as in Fig. 4 were used. The blue line in each figure represents the corresponding theoretical predictions computed from the synchronized solution (4.1) in the CL (1.8). The agreement between the numerical results and theoretical predictions is fine for cases (i) and (ii) in Fig. 5(a) and (b), and good for case (iii) in Fig. 5(c) although some fluctuations due to randomness are found in the latter. Thus, we observe that the synchronized solution (4.1) behaves as it is an asymptotically stable one in the CKM (1.7).
Figure 6. Maximal and minimal deviations of the steady state responses from the desired motion in the CKM (1.7) with $n=1000$, $K=0.5$, $V_{1},b_{0}=1$ and $V_{0}=1$ when the feedback gain $b_{1}$ is changed: (a) $(a,p)=(1,1)$ in case (i); (b) $(0.5,0.5)$ in case (ii); (c) $(a,p,\gamma)=(0.5,0.5,0.3)$ in case (iii). Here $\max_{i\in[n]}(u_{i}^{n}(t)-V(t))$ and $\min_{i\in[n]}(u_{i}^{n}(t)-V(t))$, $i\in[n]$, for $t>0$ sufficiently large are plotted as red dots. The blue line represents the corresponding theoretical predictions ±$\pm\Delta u$ computed from the synchronized solution (4.1) in the CL (1.8), where $\Delta u$ is given by (6.3).
Figures 6(a), (b) and (c) display the maximal and minimal deviations of the steady state responses from the desired motion in cases (i), (ii) and (iii), respectively, when the feedback gain $b_{1}$ is varied. The blue line in each figure represents the corresponding theoretical predictions ±$\pm\Delta u$ with
$\Delta u=\arcsin\left(\frac{a}{2(pKC+b_{1})}\right)$ (6.3)
computed from the synchronized solution (4.1) in the CL (1.8), where $C$ satisfies (4.2), and the green line the critical value of $b_{1}$ given by (4.4) at which the solution suddenly appears when $b_{1}$ is increased. We observe that the steady state responses tend to the desired motion as $b_{1}\to\infty$. The agreement between the numerical results and theoretical predictions is fine for cases (i) and (ii) in Fig. 6(a) and (b), and good for case (iii) in Fig. 6(c), again.
Acknowledgments
This work was partially supported by JSPS KAKENHI Grant Number JP23K22409. The author thanks Donggeon Kim for his assistance in numerical computations.
Appendix A Derivation of (1.9) and (1.13)
We first derive the solution (1.9) to the CL (1.8). We substitute $u(t,x)=\tilde{U}(x)+V(t)$ into (1.8) to obtain
$V_{1}=$ $\omega(x)+pK\cos\tilde{U}(x)\int_{I}\sin\tilde{U}(y)\mathrm{d}y$
$-pK\sin\tilde{U}(x)\int_{I}\cos\tilde{U}(y)\mathrm{d}y-b_{1}\sin\tilde{U}(x)+b_{0}$
$=$ $\omega(x)-\tilde{C}\sin(\tilde{U}(x)-\psi)+b_{0},$
where
$\tilde{C}\cos\psi=pK\int_{I}\cos\tilde{U}(y)\mathrm{d}y+b_{1},\quad\tilde{C}\sin\psi=pK\int_{I}\sin\tilde{U}(y)\mathrm{d}y.$ (A.1)
Obviously, $\tilde{C}\neq 0$ since $\omega(x)-V_{1}+b_{0}\equiv 0$ otherwise. Hence, we have
$\sin\left(\tilde{U}(x)-\psi\right)=\frac{\omega(x)-V_{1}+b_{0}}{\tilde{C}},$ (A.2)
which yields
$\tilde{U}(x)=\arcsin\left(\frac{\omega(x)-V_{1}+b_{0}}{\tilde{C}}\right)+\psi$ (A.3)
and
$\begin{split}\tilde{C}\cos\psi=&pK\cos\psi\int_{I}\sqrt{1-\left(\frac{\omega(y)-V_{1}+b_{0}}{\tilde{C}}\right)^{2}}\mathrm{d}y+b_{1},\\ \tilde{C}\sin\psi=&pK\sin\psi\int_{I}\sqrt{1-\left(\frac{\omega(y)-V_{1}+b_{0}}{\tilde{C}}\right)^{2}}\mathrm{d}y\end{split}$ (A.4)
by (A.1) since by (1.11)
$\int_{I}(\omega(y)-V_{1}+b_{0})\mathrm{d}y=0.$
From (A.4) we obtain $\psi=0$ and
$\tilde{C}=pK\int_{I}\sqrt{1-\left(\frac{\omega(y)-V_{1}+b_{0}}{\tilde{C}}\right)^{2}}\mathrm{d}y+b_{1}.$ (A.5)
Comparing (A.5) with (1.10), we see that $\tilde{C}=pKC+b_{1}$, so that $\tilde{U}(x)=U(x)$ by (A.3). Thus, we obtain the solution (1.9).
We next derive the solution (1.13) to the CKM (1.7) with $w_{ij}^{n}=p$, $i,j\in[n]$, and $\alpha_{n}=1$. We substitute $u_{i}^{n}(t)=\tilde{U}_{i}^{n}+V(t)$ into (1.7) to obtain
$V_{1}=$ $\omega_{i}^{n}+\frac{pK}{n}\cos\tilde{U}_{i}^{n}\sum_{j=1}^{n}\sin\tilde{U}_{j}^{n}-\frac{pK}{n}\sin\tilde{U}_{i}^{n}\sum_{j=1}^{n}\cos\tilde{U}_{j}^{n}-b_{1}\sin\tilde{U}_{i}^{n}+b_{0}$
$=$ $\omega_{i}^{n}-\tilde{C}_{\mathrm{D}}\sin(\tilde{U}_{i}^{n}-\psi)+b_{0}$
where
$\tilde{C}_{\mathrm{D}}\cos\psi=\frac{pK}{n}\sum_{j=1}^{n}\cos\tilde{U}_{j}^{n},+b_{1}\quad\tilde{C}_{\mathrm{D}}\sin\psi=\frac{pK}{n}\sum_{j=1}^{n}\sin\tilde{U}_{j}^{n}.$ (A.6)
Hence, we have
$\tilde{U}_{i}^{n}=\arcsin\left(\frac{\omega_{i}^{n}-V_{1}+b_{0}}{\hat{C}_{\mathrm{D}}}\right)+\psi,\quad i\in[n],$ (A.7)
and
$\begin{split}&\tilde{C}_{\mathrm{D}}\cos\psi=\frac{pK}{n}\cos\psi\sum_{j=1}^{n}\sqrt{1-\left(\frac{\omega_{j}^{n}-V_{1}+b_{0}}{\tilde{C}_{\mathrm{D}}}\right)^{2}}+b_{1},\\ &\tilde{C}_{\mathrm{D}}\sin\psi=\frac{pK}{n}\sin\psi\sum_{j=1}^{n}\sqrt{1-\left(\frac{\omega_{j}^{n}-V_{1}+b_{0}}{\tilde{C}_{\mathrm{D}}}\right)^{2}}\end{split}$ (A.8)
by (A.6) since by (1.16)
$\frac{1}{n}\sum_{j=1}^{n}\omega_{i}^{n}-V_{1}+b_{0}=0.$
From (A.8) we obtain $\psi=0$ and
$\tilde{C}_{\mathrm{D}}=\frac{pK}{n}\sum_{j=1}^{n}\sqrt{1-\left(\frac{\omega_{j}^{n}-V_{1}+b_{0}}{\tilde{C}_{\mathrm{D}}}\right)^{2}}+b_{1}.$ (A.9)
Comparing (A.9) with (1.14), we see that $\tilde{C}_{\mathrm{D}}=pKC_{\mathrm{D}}+b_{1}$, so that $\tilde{U}_{i}^{n}=U_{i}^{n}$, $i\in[n]$, by (A.7). Thus, we obtain the solution (1.13).